OpenStudy (anonymous):

Are you familiar with line integrals in vector fields? if So please help with the question i post below

OpenStudy (anonymous):

let C_ij be the boundary of the square [i,i+1]x[j,j+1], oriented clockwise, and \[F = <x^2y^2+2, 3x+y>.\] Calculate \[\sin(\sum_{i=4}^{14}\sum_{j=7}^{17} \int\limits_{C_ij}Fdotdr)\]

OpenStudy (anonymous):

where dot is the dot operator not multiplication

OpenStudy (anonymous):

and F is a vector field which should be in bold (or with an arrow on top), and dr is a vector as well

OpenStudy (anonymous):

okk this is of the form \[\int\limits_{}^{} (x^2y^2+2 )dx + \int\limits_{}^{}(3x+y)dy\] .. Now apply greens theorem to get \[\int\limits_{j}^{j+1} \int\limits_{i}^{i+1}(3 - 2x^2y)dxdy\] ... solve this and use summation to find the value.

OpenStudy (anonymous):

Great! i'll give it a try tomorrow morning thank you

OpenStudy (anonymous):

one question though, is there a way that I wouldn't have to integrate that 10 times? like do some terms automatically cancel out somehow?

OpenStudy (anonymous):

you don't need to integrate 10 times.. you need to integrate once ...Then you need to sum it 10 times (integration would give you number in terms of i and j)